Twenty percent of the paintings in a gallery are not originals. A collector buys a painting he thinks is an original. He has probability 0.10 of mistaking a fake for an original but never rejects an original as a fake, What is the (conditional) probability the painting he purchases is an original? (Use Bayes’ Theorem.)
Answer #1

P(think original)=P(original and collector thinks original)+P(not original and collector thinks original)

=(1-0.2)*1+0.2*0.1=0.82

hence (conditional) probability the painting he purchases is an original given collector think it as original =P(original and collector thinks original)/P(think original)

=(1-0.2)*1/0.82=0.975610